OpenStudy (anonymous):

How to solve |5-6t|??

OpenStudy (jdoe0001):

what's the full context? I mean, absolute value of anything is just a non-negative number of whatever you get inside it so |2-5| =3 |5-2| = 3

OpenStudy (anonymous):

my answers are t√61 √61 √11 √11t

OpenStudy (jdoe0001):

what's the full equation? |5-6t| = 0?